# HRI-RECAPP-2024-05

### v1

@@ Abstract
We show how the 2-Higgs Doublet Model (2HDM) Type-I can explain some excesses recently seen at the Large Hadron Collider (LHC) in $\gamma\gamma$ and $\tau^{+}\tau^{-}$ final states in turn matching Large Electron Positron (LEP) data in $b\bar{b}$ signatures, all anomalies residing over the 90-100 GeV or so region. The explanation to such anomalous data is found in the aforementioned scenario when in inverted mass hierarchy, in two configurations: i) when the lightest CP-even Higgs state is alone capable of reproducing the excesses; ii) when a combination of such a state and the CP-odd Higgs boson is able to do so. To test further this scenario, we present some Benchmark Points (BPs) of it amenable to phenomenological investigation.
@@ I Introduction
A long-standing anomaly existing in LEP collider data [1] is the one hinting at the possibility of $e^{+}e^{-}\to Zh$ events being produced therein, with a Higgs boson state $h$ with a mass of approximately 98 GeV decaying into ${b\bar{b}}$ pairs [2]. More recently, the CMS collaboration at the LHC has found an excess near 95 GeV in di-photon events in two separate analyses [3, 4]. In fact, they also reported an excess in $\tau^{+}\tau^{-}$ pairs, again, around a mass of about 98 GeV. Finally, ATLAS also observed an excess at around 95 GeV in di-photon events, thereby aligning with CMS although, especially when including ‘look elsewhere’ effects, their findings are far less significant than the CMS ones. Altogether, in view of the limited mass resolution of the di-jet invariant mass at LEP, this older anomaly may well be consistent with the excesses seen by CMS (and, partially) ATLAS in the $\gamma\gamma$ and, even more so, $\tau^{+}\tau^{-}$ final states (as the mass resolution herein is also rather poor).
As a consequence of this credible mass overlap, many studies [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] have tested the possibility of simultaneously fitting these excesses within Beyond the Standard Model (BSM) frameworks featuring a non-SM Higgs state lighter than 125 GeV, i.e., than the one observed at the LHC in 2012 [36, 37].
@@
In this study, we show that solutions of the same kind (i.e., both a fully CP-even and a mixed CP-even/odd one) can also be found in a 2HDM Type-I, wherein only one Higgs doublet gives mass to all SM fermions, again, satisfying the aforementioned theoretical requirements and experimental constraints.
The paper is organised as follows. In the next section, we review the theoretical framework of the 2HDM Type-I. Then we discuss the theoretical and experimental constraints applied to such a BSM scenario in our study, after which we move on to show how the latter can naturally explain the discussed anomalous data in various configurations of its parameters space. Our conclusions then follow.
@@ II The 2HDM Type-I
Among the various BSM scenarios, the 2HDM can be considered as a simple extension of the SM. The scalar sector of this model comprises two complex scalar fields $\phi_{1}$ and $\phi_{2}$ which transform as a doublet under the Electro-Weak (EW) gauge group $SU(2)_{L}\times U(1)_{Y}$ with hypercharge $Y=1$. For a detailed overview of this model interested reader can look into  [44]. The most general gauge invariant CP-conserving scalar potential can be written as
@@
Given the hermiticity of the scalar potential, all the potential parameters of Eq.(1) must be real. In order to prevent tree level Flavour Changing Neutral Currents (FCNCs), one can postulate an additional discrete $\mathcal{Z}_{2}$ symmetry under which the scalar fields transform as $\phi_{1}\to\phi_{1}$, $\phi_{2}\to-\phi_{2}$. One can realise that the term proportional to $m^{2}_{12}$, $\lambda_{6}$ and $\lambda_{7}$ in Eq.(1) violates this $\mathcal{Z}_{2}$ symmetry explicitly. Therefore the potential can be expressed in the following form:
@@
Both these Higgs fields, $\phi_{1}$ and $\phi_{2}$ acquire a non-zero vacuum expectation value (vev) (i.e., $\langle\phi_{1}\rangle=v_{1}$ and $\langle\phi_{2}\rangle=v_{2}$) and spontaneously break the EW gauge symmetry down to $U(1)_{\rm EM}$. After the symmetry breaking, the $W$ and $Z$ boson becomes massive and the scalar sector contains five physical Higgs bosons - two CP-even $\{H,h\}$, one CP-odd $A$ and a pair of charged states $H^{\pm}$ with masses $m_{H},m_{h},m_{A}$ and $m_{H^{\pm}}$, respectively. Both these vevs $v_{1}$ and $v_{2}$ are related to the EW scale $v=\sqrt{v^{2}_{1}+v^{2}_{2}}$ = 246 GeV. The ratio between these two vevs can be parameterised as $\tan\beta=\frac{v_{2}}{v_{1}}$. In addition, the mixing angle between the CP-even states $\{H,h\}$ can be parametrised as $\alpha$. For present study we consider the inverted hierarchy between the CP-even mass eigenstates. This particular choice alters the usual interpretation of the mass spectrum and the couplings  [^1] For example, in the inverted mass hierarchy the alignment limit corresponds to $\cos(\beta-\alpha)\to 1$. However, we will not conform to this limit in the present study.. Hereafter we align $H$ as to be the SM like Higgs boson and $h$ to be the lighter scalar. In Eq. (3) we present the relations between the potential parameters $\lambda_{i}$’s with the physical parameters of the model:
@@
The scalar potential given in Eq.(2) clearly has six independent parameters, and the $m_{ii}^{2}$ terms can be traded off in terms of $\lambda_{i}$ using the extremization conditions on the potential. We find relations of the parameters of the potentials given in equation Eq.(3) in terms of physical scalar masses and the angles $\{\beta,\alpha\}$ and use those as the input parameters for further analysis.
The right handed up/down quarks and lepton fields are also charged under the aforementioned $Z_{2}$ symmetry and transform as $u^{i}_{R}\to-u^{i}_{R}$, $d^{i}_{R}\to-d^{i}_{R}$ and $\ell^{i}_{R}\to-\ell^{i}_{R}$, respectively. From these charge assignment one realises that all charged fermions exclusively couple to the $\Phi_{2}$ field, leading to the traditional 2HDM Type-I scenario. In Eq.(4) we write down the Yukawa part of the Lagrangian in the mass eigenstate basis.
@@
Here $m_{f}$ is the fermion mass, $V$ is the CKM matrix and $P_{L/R}=\frac{1\pm\gamma_{5}}{2}$ are the projection operators. The explicit form of the scaling function $\kappa_{i}$ are detailed in Table  1.
@@ III Constraints
In this section, we describe different theoretical and experimental constraints which are required to restrict the parameter space of the Type-I 2HDM.
@@ III.1 Theoretical Constraints
  • Vacuum Stability: The vacuum stability conditions ensure that the potential must be bounded from below in all possible field direction. To achieve these, the $\lambda_{i}$’s parameters must follow certain relationship such that the quartic terms in the potential must dominate in large field values. Here we list down the conditions on $\lambda_{i}$’s which is needed to meet the stability criteria, that prevents the potential from becoming infinitely negative, [45]
@@
  • Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies. At tree level, unitarity imposes specific conditions on the energy growth of all possible $2\to 2$ scattering processes. Ref[46, 47] derives the unitarity conditions for the 2HDM model explicitly. According to the unitarity constraint, the following relations should be obeyed:
@@
    $\begin{split}u_{1}&=\frac{1}{2}(\lambda_{1}+\lambda_{2}\pm\sqrt{(\lambda_{1}-\lambda_{2})^{2}+4|\lambda_{5}|^{2}}),\\ u_{2}&=\frac{1}{2}(\lambda_{1}+\lambda_{2}\pm\sqrt{(\lambda_{1}-\lambda_{2})^{2}+4\lambda_{4}^{2}}),\\ u_{3}&=\frac{1}{2}(3(\lambda_{1}+\lambda_{2})\pm\sqrt{9(\lambda_{1}-\lambda_{2})^{2}+4(2\lambda_{3}+\lambda_{4})^{2}}),\\ u_{4}&=\lambda_{3}+2\lambda_{4}\pm 3|\lambda_{5}|,\\ u_{5}&=\lambda_{3}\pm|\lambda_{5}|,\\ u_{6}&=\lambda_{3}\pm\lambda_{4}.\\ \end{split}$
  • Perturbativity: The perturbativity condition on the parameters of the scalar potential, which imposes an upper limit on all the quartic couplings, demands that, for all $i$ values, one has $\lambda_{i}\leq|4\pi|.$
@@ III.2 Experimental Constraints
  • EW Precision Tests We evaluated the EW precision constraints by computing the $S,T$ and $U$ parameters using the SPheno package [48], with the model files written in SARAH [49]. These so-called ‘oblique parameters’ provide stringent constraints on new physics, thereby demanding that any extension to the SM Higgs sector should conform to high precision data from LEP (primarily). The numerical values of these observables are [50]
    $S=-0.02\pm 0.10,\,\,\,\,\,\,\,T=0.03\pm 0.12,\,\,\,\,\,\,\,U=0.01\pm 0.11.$
  • BSM Higgs Boson Exclusion We assessed the exclusion limits from direct searches for the BSM scalars at the LHC, LEP and the Tevatron. These exclusion limits were evaluated at the $95\%$ Confidence Level (C.L.) using the HiggsBounds-6 [51] module via the HiggsTools [52] package. In our analysis we have also demanded that our lighter Higgs must comply with the results of  [53], where the scalars are produced in association with a massive vector boson or a top anti-quark pair and further decays via leptonic modes.
  • SM-Like Higgs Boson Discovery We examined the compatibility of our 125 GeV Higgs boson with the discovered SM-like Higgs boson using a goodness-of-fit test. Specifically, we calculated the $\chi$-square value with HiggsSignals-3 [54] via HiggsTools, comparing the predicted signal strengths of our Higgs boson to those observed experimentally. We retained the parameter spaces that satisfies the condition $\chi_{125}^{2}<189.42$, corresponding to a $95\%$ C.L. with $159$ degrees of freedom.
  • Flavour Physics We incorporated constraints from $B$-physics observables, which are sensitive to potential new physics contributions in loop mediated FCNC processes. Specifically, we tested the most stringent bound on the Branching Ratio ($\mathcal{BR}$) of the $B\rightarrow X_{s}\gamma$ decay using Next-to-Leading Order (NLO) calculations in QCD as discussed in [55].
@@
    where, $\mathcal{BR}_{\rm SL}$ is the semi-leptonic branching ratio and $\Gamma_{SL}$ is the semi-leptonic decay width.
    We took our input parameters from the most recent values from the Particle Data Group (PDG) compilation [50], as follows:
@@
    Other $B$-physics observables, like $\mathcal{BR}(B^{+}\rightarrow\tau^{+}\nu_{\tau})$, $\mathcal{BR}(D_{s}\rightarrow\tau\nu_{\tau})$, $\mathcal{BR}(B_{s}\rightarrow\mu^{+}\mu^{-})$ and $\mathcal{BR}(B^{0}\rightarrow\mu^{+}\mu^{-})$ have been taken care of by using the FlavorKit tool [56] provided by SPheno package [48]. Our calculated $b\rightarrow s\gamma$ results were also found to be consistent with the FlavorKit tool.
@@ IV Explaining the anomalies
The primary objective of this paper is to investigate whether the 2HDM Type-I can explain the excesses observed at the LHC and LEP data over the $92-98$ GeV or so mass range. To do so, we need to define the signal strength corresponding to these three excesses. The signal strength is formulated as a ratio between the observed number of events to the expected number of events for a hypothetical SM Higgs of mass $95$ GeV. Assuming the Narrow Width Approximation (NWA), the signal strength for the $\tau^{+}\tau^{-},\gamma\gamma$ and $b\bar{b}$ channels can be parameterised as cross section ($\sigma$) times the branching ratio ($\mathcal{BR}$),
@@
Although the ditau excess is most prominent around $100$ GeV and the $b\bar{b}$ excess near $98$ GeV, a search around $95$ GeV could provide a unified explanation for all these three anomalies. This is because the mass resolution in the ditau final states is substantially large, and the LEP excess, associated with the $b\bar{b}$ final states, is also broad. Therefore, a common origin for these excesses may plausibly reside around $95$ GeV.
In our analysis, we have combined the di-photon measurements from the ATLAS and CMS experiments, denoted as $\mu_{\gamma\gamma}^{\rm ATLAS}$ and $\mu_{\gamma\gamma}^{\rm CMS}$, respectively. The ATLAS measurement yields a central value of $0.18\pm 0.1$ [61] while the CMS measurement yields a central value of $0.33^{+0.19}_{-0.12}$ [62]. The combined measurement, denoted by $\mu_{\gamma\gamma}^{\rm ATLAS+CMS}$ is determined by taking the average of these two central values, assuming both these measurements are uncorrelated. The corresponding combined uncertainty is calculated by adding corresponding uncertainties in quadrature. To determine if the observed excess can be explained through our model or otherwise, we perform a $\chi^{2}$ analysis using the central values $\mu^{\rm exp}$ and the $1\sigma$ uncertainties $\Delta\mu^{\rm exp}$ associated with the signal related to the excesses as defined in Eq. (7). The contribution to the $\chi^{2}$ for each of the channel is calculated using the equation
@@
Hence, the resulting $\chi^{2}$ which we will use to determine if the excesses are explained by the 2HDM Type-I, or otherwise, is the following:
@@
We test this BSM scenario into two cases: firstly, we consider both the CP-even and CP-odd (i.e., $\phi=h+A$. except for $b\bar{b}$ where $\phi=h$) Higgs states simultaneously in explaining the anomaly and, secondly, we only exploit the CP-even Higgs (i.e., $\phi=h$) in order to explain it. Hence, we align our $H$ state (recall that we have $m_{h}<m_{H}$) with the SM Higgs boson, so that $m_{H}=125$ GeV, and start a Monte Carlo (MC) sampling of the various input parameters.

### v2

@@ Abstract
We show how the 2-Higgs Doublet Model (2HDM) Type-I can explain some excesses recently seen at the Large Hadron Collider (LHC) in $\gamma\gamma$ and $\tau^{+}\tau^{-}$ final states in turn matching Large Electron Positron (LEP) data in $b\bar{b}$ signatures, all anomalies residing around 95 GeV. The explanation to such anomalous data is found in the aforementioned scenario when in inverted mass hierarchy, in two configurations: i) when the lightest CP-even Higgs state is alone capable of reproducing the excesses; ii) when a combination of such a state and the CP-odd Higgs boson is able to do so. To test further this scenario, we present some Benchmark Points (BPs) of it amenable to phenomenological investigation.
@@ I Introduction
A long-standing anomaly existing in LEP collider data [1] is the one hinting at the possibility of $e^{+}e^{-}\to Zh$ events being produced therein, with a Higgs boson state $h$ with a mass of approximately 98 GeV decaying into $b\bar{b}$} pairs [2]. More recently, the CMS collaboration at the LHC has found an excess near 95 GeV in di-photon events in two separate analyses [3, 4]. In fact, they also reported an excess in $\tau^{+}\tau^{-}$ pairs, again, around a mass of about 98 GeV. Finally, ATLAS also observed an excess at around 95 GeV in di-photon events, thereby aligning with CMS, although, especially when including ‘look elsewhere’ effects, their findings are far less significant than the CMS ones. Altogether, in view of the limited mass resolution of the di-jet invariant mass at LEP, this older anomaly may well be consistent with the excesses seen by CMS (and, partially) ATLAS in the $\gamma\gamma$ and, even more so, $\tau^{+}\tau^{-}$ final states (as the mass resolution herein is also rather poor). As a consequence of this credible mass overlap, many studies [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] have tested the possibility of simultaneously fitting these excesses within Beyond the Standard Model (BSM) frameworks featuring a non-SM Higgs state lighter than 125 GeV, i.e., than the one observed at the LHC in 2012 [38, 39].
@@
In this study, we would like to analyse whether solutions of the same kind (i.e., both a fully CP-even and a mixed CP-even/odd one) can also be found in a 2HDM Type-I, wherein only one Higgs doublet gives mass to all SM fermions, again, satisfying the aforementioned theoretical requirements and experimental constraints.
The paper is organised as follows. In the next section, we briefly review the theoretical framework of the 2HDM Type-I. Then we discuss the theoretical and experimental constraints applied to such a BSM scenario in our study, which are described in section III. In section IV we show how the aforementioned configurations can naturally explain the discussed anomalous data. We then summarise our findings in section V.
@@ II The 2HDM Type-I
Among the various BSM scenarios, the 2HDM can be considered as a simple extension of the SM. In fact, the (pseudo)scalar sector of this model comprises two complex Higgs fields, $\phi_{1}$ and $\phi_{2}$, which transform as a doublet under the Electro-Weak (EW) gauge group $SU(2)_{L}\times U(1)_{Y}$ with hypercharge $Y=1$. (For a detailed overview of this model interested reader can look into  [46].) The most general gauge invariant CP-conserving scalar potential can be written as
@@
Given the hermiticity of the scalar potential, all the potential parameters of Eq. (1) must be real. In order to prevent tree level Flavour Changing Neutral Currents (FCNCs), one can postulate an additional discrete $\mathcal{Z}_{2}$ symmetry under which the scalar fields transform as $\phi_{1}\to\phi_{1}$, $\phi_{2}\to-\phi_{2}$. One then realises that the term proportional to $m^{2}_{12}$, $\lambda_{6}$ and $\lambda_{7}$ in Eq. (1) violates this $\mathcal{Z}_{2}$ symmetry explicitly. Therefore the potential can be expressed in the following form:
@@
Both these Higgs fields, $\phi_{1}$ and $\phi_{2}$, acquire a non-zero Vacuum Expectation Value (VEV) (i.e., $\langle\phi_{1}\rangle=v_{1}$ and $\langle\phi_{2}\rangle=v_{2}$) and spontaneously break the EW gauge symmetry down to $U(1)_{\rm EM}$. After symmetry breaking, the $W^{\pm}$ and $Z$ boson becomes massive and the (pseudo)scalar sector contains five physical Higgs bosons - two CP-even $\{H,h\}$, one CP-odd $A$ and a pair of charged states $H^{\pm}$ with masses $m_{H},m_{h},m_{A}$ and $m_{H^{\pm}}$, respectively. Both these VEVs $v_{1}$ and $v_{2}$ are related to the EW scale $v=\sqrt{v^{2}_{1}+v^{2}_{2}}$ = 246 GeV. The ratio between these two VEVs can be parameterised as $\tan\beta=\frac{v_{2}}{v_{1}}$. In addition, the mixing angle between the CP-even states $\{H,h\}$ can be parametrised as $\alpha$. For the present study we consider the inverted hierarchy between the CP-even mass eigenstates. This particular choice alters the usual interpretation of the mass spectrum and the couplings [^1] For example, in the inverted mass hierarchy the alignment limit corresponds to $\cos(\beta-\alpha)\to 1$. However, we will not conform strictly to this limit in the present study.. Hereafter we align $H$ as to be the SM like Higgs boson and $h$ to be the lighter scalar state. In Eq. (3) we present the relations between the potential parameters $\lambda_{i}$’s with the physical parameters of the model:
@@
The scalar potential given in Eq. (2) clearly has six independent parameters, so that the $m_{ii}^{2}$’s can be traded off in terms of $\lambda_{i}$’s using the extremisation conditions on the potential. We find relations of the parameters of the potentials given in Eq. (3) in terms of physical scalar masses and the angles $\{\beta,\alpha\}$, then use those as the input parameters for further analysis.
The right-handed up/down quarks and lepton fields are also charged under the aforementioned $Z_{2}$ symmetry and transform as $u^{i}_{R}\to-u^{i}_{R}$, $d^{i}_{R}\to-d^{i}_{R}$ and $\ell^{i}_{R}\to-\ell^{i}_{R}$, respectively. From these charge assignments one realises that all charged fermions exclusively couple to the $\Phi_{2}$ field, leading to the traditional 2HDM Type-I scenario. In Eq. (4) we write down the Yukawa part of the Lagrangian in the mass eigenstate basis.
@@
Here $m_{f}$ is the fermion mass, $V$ is the Cabibbo-Kobayashi-Maskawa (CKM) matrix and $P_{L/R}=\frac{1\pm\gamma_{5}}{2}$ are the projection operators. The explicit form of the scaling functions $\kappa_{i}$’s (also called coupling modifiers) are detailed in Table.1.
@@ III Constraints
In this section, we describe different theoretical and experimental constraints which are required to restrict the parameter space of the 2HDM Type-I.
@@ III.1 Theoretical Constraints
  • Vacuum Stability: The vacuum stability conditions ensure that the potential must be bounded from below in all possible field direction. To achieve this, the $\lambda_{i}$ parameters must follow certain relationships such that the quartic terms in the potential must dominate for large field values. Here we list down the conditions on the $\lambda_{i}$’s which are needed to meet the stability criteria, which prevents the potential from becoming infinitely negative, [47]
@@
  • Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies. At tree level, unitarity imposes specific conditions on the energy growth of all possible $2\to 2$ scattering processes. Ref. [48, 49] derives the unitarity conditions for the 2HDM model explicitly. According to the unitarity constraint, the following relations should be obeyed:
@@
[…]
  • Perturbativity: The perturbativity condition on the parameters of the Higgs potential, which imposes an upper limit on all the quartic couplings, demands that, for all $i$ values, one has $\lambda_{i}\leq|4\pi|.$
@@ III.2 Experimental Constraints
  • EW Precision Tests We evaluated the EW precision constraints by computing the $S,T$ and $U$ parameters using the SPheno package [50], with the model files written in SARAH [51]. These so-called ‘oblique parameters’ provide stringent constraints on new physics masses and relevant couplings. Therefore any BSM theory should conform to these precision data which are primarily collected by LEP, SLC and Tevatron. The numerical values, with correlation coefficients of +0.92+0.92 between S and T plus 0.68-0.68 (0.87-0.87) between S and U (T and U) are [52]
    $S=0.04\pm 0.11,\,\,\,\,\,\,\,T=0.09\pm 0.14,\,\,\,\,\,\,\,U=-0.02\pm 0.11.$
  • BSM Higgs Boson Exclusions We assessed the exclusion limits from direct searches for BSM (pseudo)scalars states at the LEP, Tevatron and the LHC. These exclusion limits were evaluated at the $95\%$ Confidence Level (C.L.) using the HiggsBounds-6 [53] module via the HiggsTools [54] package. In our analysis we have also demanded that our lighter Higgs must comply with the results of  [55], where the Higgs particles are produced in association with a massive vector boson or a top anti-quark pair and further decays via leptonic modes.
  • SM-Like Higgs Boson Discovery We examined the compatibility of our 125 GeV Higgs boson with the discovered SM-like Higgs boson using a goodness-of-fit test. Specifically, we calculated the $\chi$-square value with HiggsSignals-3 [56] via HiggsTools, comparing the predicted signal strengths of our Higgs boson to those observed experimentally. We retained the parameter spaces that satisfies the condition $\chi_{125}^{2}<189.42$ [45, 44], corresponding to a $95\%$ C.L. with $159$ degrees of freedom. The choice of the upper bound on $\chi^{2}<189.42$, corresponds to a global $\chi^{2}$ goodness of fit test at $95\%$ C.L. ensuring that the predicted properties of the heavier CP even Higgs state (with the mass $m_{H}=125$ GeV) of the underlying 2HDM Type-I closely match those of the Higgs boson observed at the LHC, thereby keeping the model compatible under current experimental constraints.
  • Flavour Physics We incorporated constraints from $B$-physics observables, which are sensitive to potential new physics contributions in loop mediated FCNC processes. Specifically, we tested the most stringent bound on the Branching Ratio ($\mathcal{BR}$) of the $B\rightarrow X_{s}\gamma$ decay using Next-to-Leading Order (NLO) calculations in QCD as discussed in [57]:
@@
[…]
    We took our input parameters from the most recent Particle Data Group (PDG) compilation [58], as follows:
@@
    Other $B$-physics observables, like $\mathcal{BR}(B^{+}\rightarrow\tau^{+}\nu_{\tau})$, $\mathcal{BR}(D_{s}\rightarrow\tau\nu_{\tau})$, $\mathcal{BR}(B_{s}\rightarrow\mu^{+}\mu^{-})$ and $\mathcal{BR}(B^{0}\rightarrow\mu^{+}\mu^{-})$ have been taken care of by using the FlavorKit tool [59] provided by the SPheno package [50]. Our calculated $b\rightarrow s\gamma$ results were also found to be consistent with the FlavorKit tool.
@@ IV Explaining the anomalies
The primary objective of this paper is to investigate whether the 2HDM Type-I can explain the excesses observed at the LHC and in LEP data over the 949694-96 GeV mass range. To do so, we need to define the signal strength corresponding to these three excesses. The signal strength is formulated as a ratio between the observed number of events to the expected number of events for a hypothetical SM Higgs state of mass $95$ GeV. Assuming the Narrow Width Approximation (NWA), the signal strength for the $\tau^{+}\tau^{-},\gamma\gamma$ and $b\bar{b}$ channels can be parameterised as cross section ($\sigma$) times $\mathcal{BR}$, as follows:
@@
Although the ditau excess is most prominent around $100$ GeV and the $b\bar{b}$ excess near $98$ GeV, a search around $95$ GeV could provide a unified explanation for all these three anomalies. This is because the mass resolution in the ditau final states is substantially large and the LEP excess, associated with the $b\bar{b}$ final states, is also rather broad: therefore, a common origin for these excesses may plausibly reside around $95$ GeV [20].
In our analysis, we have combined the di-photon measurements from the ATLAS and CMS experiments, denoted as $\mu_{\gamma\gamma}^{\rm ATLAS}$ and $\mu_{\gamma\gamma}^{\rm CMS}$, respectively. The ATLAS measurement yields a central value of $0.18\pm 0.1$ [64] while the CMS measurement yields a central value of $0.33^{+0.19}_{-0.12}$ [65]. The combined measurement, denoted by $\mu_{\gamma\gamma}^{\rm ATLAS+CMS}$ is determined by taking the average of these two central values, assuming these measurements are uncorrelated. The corresponding combined uncertainty is calculated by adding the individual uncertainties in quadrature. To determine whether the observed excess can be explained through our model or otherwise, we perform a $\chi^{2}$ analysis using the central values $\mu^{\rm exp}$ and the $1\sigma$ uncertainties $\Delta\mu^{\rm exp}$ associated with the signal related to the excesses as defined in Eq. (7). The contribution to the $\chi^{2}$ for each of the channel is calculated using the equation
@@
Hence, the resulting $\chi^{2}$, which we will use to determine if the excesses can be achieved by the allowed region of the parameter space of 2HDM Type-I, or otherwise, is the following:
@@
We test this BSM scenario for two cases: firstly, we consider both the CP-even and CP-odd Higgs states simultaneously (i.e., $\phi=h+A$, except for $b\bar{b}$ where $\phi=h$ as the CP-odd scalar $A$ does not couple to $ZZ$ mode) in explaining the anomaly and, secondly, we only exploit the CP-even Higgs state (i.e., $\phi=h$) in order to explain it. Hence, we align our $H$ state (recall that we have $m_{h}<m_{H}$) with the SM Higgs boson, so that $m_{H}=125$ GeV, and start a Monte Carlo (MC) sampling of the various input parameters.